OpenStudy (aravindg):

solve for x and y ax+hy+g=0 hx+by+f=0

OpenStudy (aravindg):

help @UnkleRhaukus , @saifoo.khan , @cwrw238

OpenStudy (aravindg):

@kropot72

OpenStudy (kropot72):

ax + hy + g = 0 Dividing both sides by a gives: x + (h/a)y + g/a = 0 .....................(1) hx + by + f = 0 Dividing both sides by h gives: x + (b/h)y + f/h = 0 .....................(2) Subtracting (2) from (1) gives: (h/a)y - (b/h)y + g/a - f/h = 0 y(h/a - b/h) = f/h - g/a y = (f/h - g/a) / (h/a - b/h)

OpenStudy (aravindg):

should i do the same process again to get x?@kropot72

OpenStudy (kropot72):

@AravindG Yes, I recommend that you use the same process to find x. Substitution of the value for y looks to be rather more complicated.

OpenStudy (aravindg):

ya tht was what i was also thinking..thanks a lot!!!!!!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=Solve [a+x+%2B++h+y+%2B+g+%3D%3D+0+%26%26+h+x+%2B+b+y+%2B+f+%3D%3D+0%2C+{x%2C+y}]

OpenStudy (anonymous):

Copy and paste in your browser the above URL